OpenStudy (anonymous):

The ratio of the numerator to the denominator of a certain fraction is one to four. If three is added to the numerator and subtracted from the denominator, the new fraction reduces to one-third. What is the original fraction? a) 12/45 b) 15/45 c) 12/48

OpenStudy (e.mccormick):

How far did you get in setting this up?

OpenStudy (anonymous):

idk where to even start

OpenStudy (e.mccormick):

Well, you have two equations described. Did you try writing them out?

OpenStudy (e.mccormick):

\(?=\frac{1}{4}\) \(?=\frac{1}{3}\) What would the ? parts be?

OpenStudy (e.mccormick):

Have you done equations with multiple unknowns? x+y=5 2x-y=8 Things like that? Because this can be done just as easily. It is just a matter of setting up the left side of the = properly.